# High CO/H2 ratios supports an exocometary origin for a CO-rich debris disk

### v2

@@ 2.1 CRIRES+ Observations
Using CRIRES+, the high-resolution near infrared spectrograph on the VLT (Dorn et al., 2023), we observed HD 110058 and HD 131488 over two nights each, as detailed in Table 2. We observed in the K-band with the K2148 filter, which has non-contiguous wavelength orders from 1904 nm to 2452 nm. Our observation dates were chosen such that the H2 v=1-0 S(0) rovibrational absorption line at 2223 nm was shifted away from nearby weak tellurics, mainly arising from atmospheric water as can be seen in Figure 6. Our observation windows also covered the weaker H2 v=1-0 Q(1) transition at 2406.6 nm. If detected, this transition could allow us to derive the excitation temperature of the H2 gas. Also within the range of the data are the v=2-0 bands from 12CO, 13CO, and C18O. The latter 2 isotopologues will be examined in upcoming work. No attempt was made to shift CO lines in this data away from nearby tellurics, as can be seen in Figure 7.
Table 2: Overview of VLT-CRIRES+ observations of HD 110058 and HD 131488.111Note: Exposure time is expressed as number of detector integrations (NDIT) $\times$ detector integration time (DIT).
@@ 2.2 CRIRES+ Data Reduction and Calibration
We used ESOREX and the CR2RES pipeline recipes detailed in the ESO CRIRES+ pipeline user manual (version 1.3.0) to reduce the raw detector images. The standard reduction for nodding observations shown in Figure 4.4 of this manual was the basis of our reduction. High-SNR flats for CRIRES+ are taken once a month, and were used instead of the default supplied flats to improve the final SNR. We performed wavelength calibration with the Uranium-Neon lamp and Fabry-Pérot etalon. After correcting for detector non-linearity, dark current, and then carrying out flat fielding and wavelength calibration, we extracted a 1D spectrum for each of the two nodding positions. A subset of the extracted data (with the two nodding positions combined) covering the wavelength of the H2 line of interest is shown in Figure 1. This data is yet to be corrected for tellurics or blaze; as such it has arbitrary flux units and displays an upward trend with wavelength.
@@
Figure 1: Reduced and extracted 1D spectra (black lines) for a portion of the detector-order containing our line of interest for HD 131488 on 11/05/23 (top) and HD 110058 on 25/03/23 (bottom). The spectra are not flux calibrated and as such the absolute values of the y-axis in units of analogue-to-digital units (adu) are not astrophysically meaningful. The vertical green line denotes the expected location of the H2 transition of interest. The spectra are in the observatory frame, but the expected H22 location has been shifted to the rest frame of the star.
@@
Our planned analysis required achieving a high-SNR and modelling the line-to-continuum ratio in the continuum normalised spectrum as detailed in Section 3, which made a standard star observation unnecessary. We performed telluric correction using the molecfit software, version 4.3.1 (Smette et al., 2015) which has been shown to perform as well or better than corrections using standard stars on CRIRES data (Ulmer-Moll et al., 2019). We performed telluric correction and telluric-based wavelength calibration on subsets of the data around lines of interest. For more detail see Appendix A and B. Figures 6 and 7 show example spectra with telluric fits with the most prominent, yet relatively weak tellurics arising from atmospheric water, while corrected spectra are shown in Figures 2, 3, and 4. We note that, by design, the H2 line under investigation does not overlap with any telluric lines, thus the removal of tellurics does not significantly affect the spectrum at or near the line. Telluric removal was also carried out for each nodding position/observing day combination for the 12CO lines in spectral regions covering the v=2-0 rovibrational band from 2318.9 to 2333.9 nm and 2335.455 to 2350.0 nm and for the H2 v=1-0 S(0) line from 2220 to 2224.5 nm. The nodding positions were modelled separately to allow for improved wavelength calibration described in Appendix B. An advantage of using molecfit is that in addition to telluric removal, it also returns a spectral resolution for the observations, specific to each detector-order, nodding position and observing day combination.
@@
To identify the $\text{H}_{2}$ and CO lines at their rest frequencies, the reduced, telluric free spectra for each observing night were Doppler shifted to the rest frame of the star using the systemic and barycentric velocities. The barycentric velocity correction is calculated using Astropy’s SkyCoord class (Collaboration et al., 2022), whereas radial (systemic) velocities for each system were determined through modelling12CO in the CRIRES+ data (see Section 3 and Table 1). The spectra molecfit returns are telluric-corrected but not continuum-normalised. We divide the telluric-corrected spectrum with a running median with a large window size to produce continuum normalised spectra.
@@ 3 Results and Modelling
The final normalised spectra for each star are displayed in Figures 2, 3, and 4, with wavelengths in the rest frame of the star. For $\text{H}_{2}$, if detectable, we expect circumstellar gas to produce an absorption line at the v=1-0 S(0) transition wavelength of 2223.29 nm (from the HITRAN database, Gordon et al., 2022). No statistically significant detection of H2 was found for either system. Additionally, as expected for cold $\text{H}_{2}$ where the ground state should be the most populated energy level, the weaker 2406.6 nm absorption line arising from the H2 v=1-0 Q(1) transition was not detected.
For both disks, we detect 12CO absorption at a radial velocity consistent with that observed in the UV for HD 110058 by Brennan et al. (2025). The radial velocity for HD 131488 differs from the one reported in Brennan et al. (2025), but it is consistent with the CO radial velocity observed in emission with ALMA by Mac Manamon (in press) as part of the ARKS program. For HD 110058, we detect the v=2-0, J=15-14 to J=1-0 and for HD 131488 we detection the v=2-0, J=8-7 to J=1-0 transitions. More CO is present in HD 110058, and thus the absorption lines are deeper, reflected in the axis scaling of Figures 3 and 4. Additionally, the excitation temperature for HD 110058 is higher, allowing more high-J transitions to be populated.
@@ 3.1 Modelling the Reduced CRIRES+ Spectra
We modelled H2 absorption in the normalised spectra using RADIS (Pannier & Laux, 2019). RADIS relies on the HITRAN line list (Gordon et al., 2022) obtained using the HITRAN Application Programming Interface (Kochanov et al., 2016), and partition functions described in Gamache et al. (2021). Given the non-detection, for H2 we assumed the gas is in LTE and that it can be approximated as having a single column density and temperature along the line of sight to the star. On the other hand, for CO we found LTE not to be a sufficiently good description of the data and we therefore adopted a simplified non-LTE with a different kinetic temperature (setting the intrinsic line width through Doppler broadening) from the excitation temperature (setting the ratio of the populations of the various rotational levels, and therefore the relative strength of our detected lines). This is consistent with the findings and the modelling of UV rovibronic lines in the same systems (Brennan et al., 2025). Although we cannot determine whether H2 is in LTE without more information about the collider densities, the H2 (v=1-0 S(0)) transition has an Einstein A of $2.524\times 10^{-7}\text{s}^{-1}$ while the CO 12CO $v=2\rightarrow 0$ transitions have Einstein A coefficients on the order of $10^{-1}\text{s}^{-1}$. Meanwhile the collisional rate coefficients found in Thi et al. (2013); Quéméner & Balakrishnan (2009) are of comparable orders of magnitude. This means the H2 transition has a lower critical density and is therefore more likely to be in LTE than the detected CO.
@@
We assessed the goodness of fit for each pair of model parameters with a log Gaussian likelihood ($\ln{\mathcal{L}}$) function:
$\ln{\mathcal{L}}=-\frac{1}{2}\sum_{i=0}^{n}\frac{(R_{i})^{2}}{\sigma_{i}^{2}}+\mathrm{ln}(2\pi\sigma_{i}^{2})$ (1)
@@
We use a Markov-Chain Monte Carlo (MCMC) method implemented in python’s emcee package (Foreman-Mackey et al., 2013) to sample the posterior probability distribution. We used uniform priors for each parameter and ran the MCMC for 50000 steps with 100 walkers to ensure convergence and to obtain a representative posterior distribution.
@@ 3.2 Results
The posterior probability distributions from the MCMC fit are shown in Figures 10, 11, 12, and 13. There is clear degeneracy between the H2 temperature and column density which it is not possible to break with a non-detection. For larger temperatures, larger column densities are allowed as fewer H2 molecules are in the ground state energy level and thus fewer potential absorbers lie between the star and the observer. When marginalised over temperature, we obtain a column density upper limits. The column density posterior is close to a uniform distribution for low column densities as small amounts of H2 would all equally result in signals below the noise level of the data. Larger column densities would produce larger signals reaching detectability and thus these models are less likely to describe our observations.
@@
We present H2 column density upper limits without assuming any temperature for the gas in Table 4. However, to calculate the $\frac{\text{CO}}{\text{H}_{2}}$ ratios, we assume that the H2 shares the same temperature as the CO kinetic temperature derived from our observations, which is reasonable if CO and any hypothetical H2 are co-located. The choice of kinetic rather than excitation temperatures is conservative with regards to how much unseen H2 could be present as the colder excitation temperatures derived for CO would lead to lower column densities.
Given the H2 column density upper limit, and the CO detection, we can combine their column densities to derive a $\frac{\text{CO}}{\text{H}_{2}}$ ratio lower limit for the gas in both systems. As the column densities for both H2 and CO are correlated with the gas temperatures, we first sampled from the CO and H2 posterior column density distributions at the same (kinetic) temperature, resulting in a large number of $\frac{\text{CO}}{\text{H}_{2}}$ samples representing the $\frac{\text{CO}}{\text{H}_{2}}$ probability distribution given our data. We then took the $99.7^{\text{th}}$ percentile of this distribution as the $3\sigma$ lower limit for each system, reported in Tables 3 and 4.
Figure 3: The top figure shows the HD 110058 12CO data (black solid line) overlaid with the best-fit model (blue dotted line). The bottom figure shows the 12CO residuals (black solid line) and the rolling variance (green dashed line).
Figure 4: The top figure shows the HD 131488 12CO data (black solid line) overlaid with the best-fit model (blue dotted line). The bottom figure shows the 12CO residuals (black solid line) and the rolling variance (green dashed line).
Table 3: CRIRES+ H2 data forward modelling results.
HD 131488 HD 110058
log(NCO (cm-2)) $18.05^{+0.06}_{-0.04}$ $19.97^{+0.27}_{-0.20}$
@@
CO/H2 Ratio $3.09\times 10^{-5}$ $1.35\times 10^{-3}$
@@ 4.1 CO/H2 ratios across planet formation: CO depletion to enhancement from protoplanetary to debris disks
We compare our lower limit for the $\frac{\text{CO}}{\text{H}_{2}}$ ratio in the exocometary belts’ gas to values derived for protoplanetary disks listed in Table 4 and plotted in Figure 5. Protoplanetary ratios typically range from $10^{-4}$ to $10^{-6}$ and tend to drop over time within the first 10 Myr of a star’s lifetime (Favre et al., 2013; Kama et al., 2016; McClure et al., 2016; Schwarz et al., 2016; Trapman et al., 2017; Zhang et al., 2017; Bergin & Williams, 2018; Zhang et al., 2020). This is in contrast with our first direct measurements (lower limits) for the exocometary belt of HD 110058 which has markedly elevated $\frac{\text{CO}}{\text{H}_{2}}$ indicating a much more H2-poor environment compared with primordial gas in protoplanetary disks (Bergin & Williams, 2018; Zhang et al., 2020) . The conservative lower limit we obtained via modelling the H2 non-detections in the CRIRES+ K-band spectra of HD 110058 is a factor of 13 larger than the canonical $\frac{\text{CO}}{\text{H}_{2}}$ ratio of $10^{-4}$ typical of the ISM, and higher than the lower limit on $\beta$ Pictoris. For our other target HD 131488, the lower limit still formally allows (albeit at low probability) a primordial $\frac{\text{CO}}{\text{H}_{2}}$ ratio similar to that of the ISM and protoplanetary disks.
Figure 5: Ratio of CO and H2 abundances for protoplanetary disks, $\beta$ Pictoris, and the two exocometary belts in this study. Details of each point can be found in the references of Table 4. This plot has been adapted from and expanded upon using the similar plots in Bergin & Williams (2018); Zhang et al. (2020).

### v3

@@ 2.1 CRIRES+ observations
Using CRIRES+, the high-resolution near infrared spectrograph on the VLT (Dorn et al. 2023), we observed HD 110058 and HD 131488 over two nights each, as detailed in Table 2. We observed in the K-band with the K2148 filter, which has non-contiguous wavelength orders from 1904 nm to 2452 nm. Our observation dates were chosen such that the H2 v=1-0 S(0) rovibrational absorption line at 2223 nm was shifted away from nearby weak tellurics, mainly arising from atmospheric water as can be seen in Figure 6. Our observation windows also covered the weaker H2 v=1-0 Q(1) transition at 2406.6 nm. If detected, this transition could allow us to derive the excitation temperature of the H2 gas. Also within the range of the data are the v=2-0 bands from 12CO, 13CO, and C18O. The latter 2 isotopologues will be examined in an upcoming work.
Table 2: Overview of VLT-CRIRES+ observations of HD 110058 and HD 131488.111Exposure time is expressed as number of detector integrations (NDIT) $\times$ detector integration time (DIT).
@@ 2.2 CRIRES+ data reduction and calibration
We used ESOREX and the CR2RES pipeline recipes detailed in the ESO CRIRES+ pipeline user manual (version 1.3.0) to reduce the raw detector images. The standard reduction for nodding observations shown in Figure 4.4 of this manual was the basis of our reduction. High-SNR flats for CRIRES+ are taken once a month, and were used instead of the default supplied flats to improve the final SNR. We performed wavelength calibration with the Uranium-Neon lamp and Fabry-Pérot etalon. After correcting for detector non-linearity, dark current, and then carrying out flat fielding and wavelength calibration, we extracted a 1D spectrum for each of the two nodding positions. The nodding positions both contained the star but at different locations in the detector to account for localised detector systematics. A subset of the extracted data (with the two nodding positions combined) covering the wavelength of the H2 line of interest is shown in Figure 1. This data is yet to be corrected for tellurics or blaze; as such it has arbitrary flux units and displays an upward trend with wavelength.
@@
Figure 1: Reduced and extracted 1D spectra (black lines) for a portion of the detector-order containing our H2 line of interest for HD 131488 on 11/05/23 (top) and HD 110058 on 25/03/23 (bottom). The spectra are not flux calibrated and as such the absolute values of the y-axis in units of analogue-to-digital units (adu) are not astrophysically meaningful. The spectra are not blaze corrected and thus the flux tends to rise with wavelength. The vertical green line denotes the expected location of the H2 transition of interest. The spectra are in the observatory frame and the H2 location has been shifted to account for the Doppler shift of the star compared to the observatory frame.
@@
Our planned analysis required achieving a high-SNR and modelling the line-to-continuum ratio in the continuum normalised spectrum as detailed in Section 3. Since flux calibration is not necessary for this goal, a standard star observation was not needed. We performed telluric correction using the molecfit software, version 4.3.1 (Smette et al. 2015) which has been shown to perform as well or better than corrections using standard stars on CRIRES data (Ulmer-Moll et al. 2019). We performed telluric correction and additional telluric-based wavelength calibration on subsets of the data around lines of interest. For more details on the wavelength calibration and why it was useful for our data, see Appendices A and B. Figures 6 and 7 show example spectra with telluric fits with the most prominent, yet relatively weak tellurics arising from atmospheric water, while corrected spectra are shown in Figures 2, 3, and 4. We note that by design, the H2 line under investigation does not overlap with any telluric lines, thus removing tellurics does not significantly affect the spectrum at or near the line. Telluric removal was also carried out for each nodding position/observing day combination for the 12CO lines in spectral regions covering the v=2-0 rovibrational band from 2318.9 to 2333.9 nm and 2335.455 to 2350.0 nm and for the H2 v=1-0 S(0) line from 2220 to 2224.5 nm. The nodding positions were modelled separately to allow for improved wavelength calibration described in Appendix B. An advantage of using molecfit is that in addition to telluric removal, it also returns a spectral resolution for the observations, specific to each detector-order, nodding position and observing day combination.
@@
To identify the $\text{H}_{2}$ and CO lines at their rest frequencies, the reduced telluric free spectra for each observing night were Doppler shifted to barycentric frame and then to the rest frame of the star using the system radial velocities. The barycentric velocity correction is calculated using Astropy’s SkyCoord class (Astropy Collaboration et al. 2022), whereas radial (systemic) velocities for each system were determined through modelling 12CO in the CRIRES+ data (see Section 3 and Table 1). The spectra molecfit returns are telluric-corrected but not continuum-normalised. We divide the telluric-corrected spectrum with a running median with a large window size to produce continuum normalised spectra.
@@ 3 Results and modelling
The final normalised spectra for each star are displayed in Figures 2, 3, and 4, with wavelengths in the rest frame of the star. For $\text{H}_{2}$, if detectable, we expect circumstellar gas to produce an absorption line at the v=1-0 S(0) transition wavelength of 2223.29 nm (from the HITRAN database, Gordon et al. 2022). No statistically significant detection of H2 was found for either system. Additionally, as expected for cold $\text{H}_{2}$ where the ground state should be the most populated energy level, the weaker 2406.6 nm absorption line arising from the H2 v=1-0 Q(1) transition was not detected.
We detect 12CO absorption at a radial velocity consistent with that observed in the UV for HD 110058 by Brennan et al. (2024). The radial velocity for HD 131488 differs from the one reported in Brennan et al. (2024) (they report a radial velocity of $4.8^{+0.1}_{-0.1}$ km s-1 whereas we find the radial velocity to be $5.97^{+0.05}_{-0.05}$ km s-1), but it is consistent with the CO radial velocity observed in emission with ALMA by Mac Manamon et al. (2026) as part of the ARKS program. For HD 110058, we detect the v=2-0, J=15-14 to J=1-0 and for HD 131488 we detection the v=2-0, J=8-7 to J=1-0 transitions. More CO is present in HD 110058, and thus the absorption lines are deeper, reflected in the axis scaling of Figures 3 and 4. Additionally, the excitation temperature for HD 110058 is higher, allowing more high-J transitions to be populated.
@@ 3.1 Modelling the reduced CRIRES+ spectra
We modelled H2 absorption in the normalised spectra using RADIS (Pannier and Laux 2019). RADIS relies on the HITRAN line list (Gordon et al. 2022) obtained using the HITRAN Application Programming Interface (Kochanov et al. 2016), and partition functions described in Gamache et al. (2021). Given the non-detection, for H2 we assumed the gas is in LTE and that it can be approximated as having a single column density and temperature along the line of sight to the star. On the other hand, for CO we found LTE not to be a sufficiently good description of the data and we therefore adopted a simplified non-LTE with a different kinetic temperature (setting the intrinsic line width through Doppler broadening) from the excitation temperature (setting the ratio of the populations of the various rotational/vibrational levels via Boltzmann distributions, and therefore the relative strength of our detected lines). This is consistent with the findings and the modelling of UV rovibronic lines in the same systems (Brennan et al. 2024). Although we cannot determine whether H2 is in LTE without more information about the collider densities, the H2 (v=1-0 S(0)) transition has an Einstein A of $2.524\times 10^{-7}\text{s}^{-1}$ while the CO 12CO $v=2\rightarrow 0$ transitions have Einstein A coefficients on the order of $10^{-1}\text{s}^{-1}$. Meanwhile the collisional rate coefficients found in Thi et al. (2013); Quéméner and Balakrishnan (2009) are of comparable orders of magnitude. This means the H2 transition has a lower critical density and is therefore more likely to be in LTE than the detected CO.
@@
We assessed the goodness of fit for each set of model parameters with a log Gaussian likelihood ($\ln{\mathcal{L}}$) function:
$\ln{\mathcal{L}}=-\frac{1}{2}\sum_{i=0}^{n}\left(\frac{R_{i}^{2}}{\sigma_{i}^{2}}+\mathrm{ln}(2\pi\sigma_{i}^{2})\right)$ (1)
@@
We use a Markov-Chain Monte Carlo (MCMC) method implemented in python’s emcee package (Foreman-Mackey et al. 2013) to sample the posterior probability distribution. We used uniform priors given in Table 4 for each parameter.
@@ 3.2 Results
The posterior probability distributions from the MCMC fit are shown in Figures 10, 11, 12, and 13. There is clear degeneracy between the H2 temperature and column density which cannot be broken with a non-detection. For larger temperatures, larger column densities are allowed as Doppler broadening causes broader, shallower lines which are harder to detect. When marginalised over temperature, we obtain column density upper limits. The column density posterior is close to a uniform distribution for low column densities as small amounts of H2 would all equally result in signals below the noise level of the data. Larger column densities would produce larger signals reaching detectability and thus these models are less likely to describe our observations.
@@
To calculate the $\frac{\text{CO}}{\text{H}_{2}}$ ratio lower limits, we assume that the H2 shares the same temperature as the CO kinetic temperature derived from our observations (as described below), which is reasonable if CO and any hypothetical H2 are co-located. The choice of kinetic rather than excitation temperatures is conservative with regards to how much unseen H2 could be present as the colder excitation temperatures derived for CO would lead to lower column densities.
Given the H2 column density upper limit, and the CO detection, we can combine their column densities to derive a $\frac{\text{CO}}{\text{H}_{2}}$ ratio lower limit for the gas in both systems. As the column densities for both H2 and CO are correlated with the gas temperatures, we paired each CO column density sample starting from the highest log posterior probability sample with a H2 column density at the closest kinetic temperature. This gave us a representative $\frac{\text{CO}}{\text{H}_{2}}$ probability distribution accounting for the temperature correlations. This is comparable to if we had fit CO and H2 simultaneously and enforced both having the same temperature. We then took the $99.7^{\text{th}}$ percentile of this distribution as the $3\sigma$ lower limit for each system, reported in Table 3.
Figure 3: The best fit (highest log posterior probability) 12CO models, normalised HD 110058 data, and residuals. The top figure shows the HD 110058 12CO data (black solid line) overlaid with the best-fit model (blue dotted line). The bottom figure shows the 12CO residuals (black solid line) and $2\sigma$ calculated to be twice the rolling standard deviation (green dashed line).
Figure 4: The best fit (highest log posterior probability) 12CO models, normalised HD 131488 data, and residuals. The top figure shows the HD 131488 12CO data (black solid line) overlaid with the best-fit model (blue dotted line). The bottom figure shows the 12CO residuals (black solid line) and $2\sigma$ calculated to be twice the rolling standard deviation (green dashed line).
Table 3: CRIRES+ CO/H2 modelling results. [^2] Best-fit values correspond to the highest log posterior probability with $1\sigma$ uncertainties ($16^{\mathrm{th}}$$84^{\mathrm{th}}$ percentiles of the posterior distributions). The H2 column density (upper limits at the $99.7^{\text{th}}$ percentile) and CO/H2 ratios assume the H2 shares the same kinetic temperature as the CO.
Parameter HD 110058 HD 131488
log(NCO (cm-2)) $20.0^{+0.3}_{-0.2}$ $18.05^{+0.06}_{-0.04}$
vr (km s-1) $12.53^{+0.01}_{-0.01}$ $5.87^{+0.05}_{-0.05}$
@@
[…]
CO/H2 Ratio $1.35\times 10^{-3}$ $3.09\times 10^{-5}$
@@ 4.1 CO/H2 ratios across planet formation: CO depletion to enhancement from protoplanetary to debris disks
We compare our lower limit for the $\frac{\text{CO}}{\text{H}_{2}}$ ratio in the exocometary belts’ gas to values derived for protoplanetary disks listed in Table 5 and plotted in Figure 5. Protoplanetary ratios typically range from $10^{-4}$ to $10^{-6}$ and tend to drop over time within the first 10 Myr of a star’s lifetime (Favre et al. 2013; Kama et al. 2016; McClure et al. 2016; Schwarz et al. 2016; Trapman et al. 2017; Zhang et al. 2017; Bergin and Williams 2017; Zhang et al. 2020). This is in contrast with our first direct measurements (lower limits) for the exocometary belt of HD 110058 which has markedly elevated $\frac{\text{CO}}{\text{H}_{2}}$ indicating a much more H2-poor environment compared with primordial gas in protoplanetary disks (Bergin and Williams 2017; Zhang et al. 2020) . The conservative lower limit we obtained via modelling the H2 non-detections in the CRIRES+ K-band spectra of HD 110058 is a factor of 13 larger than the canonical $\frac{\text{CO}}{\text{H}_{2}}$ ratio of $10^{-4}$ typical of the ISM. For our other target HD 131488, the lower limit still formally allows a primordial $\frac{\text{CO}}{\text{H}_{2}}$ ratio similar to that of the ISM and protoplanetary disks. The degree to which an ISM-like model is consistent with the data can also be seen in Figure 2 where the blue model shows the expected line depth for an ISM-like H2 abundance relative the CO column density. It’s clear that the spectrum for HD 110058 is not consistent with an ISM-like model whereas the same model for HD 131488 would be consistent with the noise level in our data.
Figure 5: Ratio of CO and H2 abundances for protoplanetary disks, $\beta$ Pictoris, and the two exocometary belts in this study. Details of each point can be found in the references of Table 5. This plot has been adapted from and expanded upon using the similar plots in Bergin and Williams (2017); Zhang et al. (2020).